OpenStudy (anonymous):

I need help solving 30+0.006p^3=150-0.12p^2

zepdrix (zepdrix):

Let's first change our decimals to fractions. Then we can multiply through by the largest denominator to get rid of all the fractions. It'll make it a lot easier to solve :D So that p^3 term... hmm, the 6 appears to be in the thousandths place. We can write it as \(\dfrac{6}{1000}\). If we do the same for the coefficient on the p^2 term, it becomes, \(\dfrac{12}{100}\). Giving us an equation of,\[\large 30+\frac{6}{1000}p^3=150-\frac{12}{100}p^2\]

zepdrix (zepdrix):

From here we'll multiply both sides by 1000 since that is our largest denominator, \[\large 30,000+6p^3=150,000-120p^2\]

OpenStudy (anonymous):

@zepdrix 120p^2+6p^3=120000 ... How do I solve p?

zepdrix (zepdrix):

Hmm this doesn't factor nicely... thinkinggggg

zepdrix (zepdrix):

@satellite73 @phi @Hero Hmm I can't quite figure this one out D: I paged some of the smarty pants on here, maybe one of these guys will have an idea.

OpenStudy (anonymous):

lol ok

zepdrix (zepdrix):

Woops, @Hero when you make the substitution, it turns the first term into \(y^{3/2}\), that's what i was getting a little confused on this one.